A per-unit tax imposed on suppliers will shift the supply curve up by the amount of the tax.


Illustrate the deadweight loss from the tax


1
Expert's answer
2021-12-07T12:06:55-0500

"P_0" and "Q_0" are the equilibrium price and quantity before the imposition of tax, respectively, as seen in the graph above. With the tax, the supply curve shifts from Supply 0Supply 0 to Supply 1.Supply 1 by the tax amount.

Because of the tax, businesses would want to supply less.

The buyer's price would increase from "P_0" to "P_1" and the seller would receive a lower price for the good from "P_0" to "P_2" . Due to this tax, producers supply less from "Q_0" to "Q_1" .

The blue traingle represents dead weight reduction, which may be calculated as:


Dead weight loss"=\frac{1}{2}\times(P_2-P_1)\times(Q_0-Q_1)"